OpenStudy (anonymous):

How would you write the given expression as an algebraic expression of x? sin(2arctan(x))

OpenStudy (anonymous):

I'm getting (2x)/sqrt(x^2+1)

OpenStudy (anonymous):

but the answer is (2x)/(x^2+1)

OpenStudy (freckles):

so you did u=arctan(x) => tan(u)=x did you draw a right triangle for this?

OpenStudy (freckles):

also sin(2u)=2sin(u)cos(u)

OpenStudy (anonymous):

yeah i drew a triangle

OpenStudy (freckles):

I think i know what you forget to do did you multiply the bottoms?

OpenStudy (anonymous):

as in like rationalize?

OpenStudy (freckles):

what did you get for sin(u) and cos(u)?

OpenStudy (anonymous):

sin(u) = x/sqrt(x^2+1) cos(u) = 1/sqrt(x^2+1)

OpenStudy (freckles):

ok great \[2\sin(u)\cos(u)=2 \cdot \frac{x}{\sqrt{x^2+1}} \cdot \frac{1}{\sqrt{x^2+1}} =\frac{2 \cdot x \cdot 1}{\sqrt{x^2+1} \cdot \sqrt{x^2+1}}\]

OpenStudy (freckles):

remember when you multiply fractions you multiply straight across on top and straight across on bottom

OpenStudy (anonymous):

ohh you have to use the double angle formula

OpenStudy (freckles):

yes because we have sin(2u)

OpenStudy (anonymous):

i skipped that entire process and just did 2sin(u)

OpenStudy (anonymous):

ok i got it thx

OpenStudy (freckles):

sin(2u) doesn't equal 2sin(u) for all u